# Reappraisal of SU(3)-flavor breaking in $B\rightarrow DP$

### v1

@@ IV Conclusions
Motivated by recently found discrepancies between experimental data and QCD factorization, we study the anatomy of a plain SU(3)${}_{F}$ expansion in $B\rightarrow DP$ decays as extracted from the data, with no further theory assumptions. We find that, although the SU(3)${}_{F}$ limit is excluded by $>5\sigma$, the data can be explained with SU(3)${}_{F}$ breaking of $\sim 20\%$i.e., the SU(3)${}_{F}$ expansion works to the expected level.
@@
As is well known, SU(3)${}_{F}$ is an approximate symmetry of QCD as the light ($u$, $d$, $s$) quark masses are not equal to each other though they (especially $u$ and $d$) are small compared to $\Lambda_{\mathrm{QCD}}$. So when predictions made by assuming SU(3)${}_{F}$ are not upheld experimentally, it often becomes important to estimate the size of SU(3)${}_{F}$ breaking before invoking new physics.
@@
As one more piece in the puzzle, our results inform future theoretical and experimental work. By considering the complete system of decays related by the SU(3)${}_{F}$ symmetry, we show unambiguously that the underlying issue is not connected to SU(3)${}_{F}$ breaking, as suggested by the ratio in Eq. (1), which is in agreement with the data. In a scenario beyond the SM (BSM), any new physics model designed to explain the deviations has therefore to respect the SU(3)${}_{F}$ structure of the SM operators, implying bounds on the hierarchy between BSM couplings to the first and second generation of quarks.
@@
It would be very interesting to complete the picture in the future by testing the isospin sum rule between $\mathcal{B}(\bar{B}_{s}^{0}\rightarrow\pi^{-}D^{+})$ and $\mathcal{B}(\bar{B}_{s}^{0}\rightarrow\pi^{0}D^{0})$, as well as the SU(3)${}_{F}$ sum rule between $\mathcal{B}(\bar{B}^{0}\rightarrow K^{-}D_{s}^{+})$ and $\mathcal{B}(\bar{B}_{s}^{0}\rightarrow\pi^{-}D^{+})$, see Eqs. (23, 24, 29). In order to achieve this, measurements of the suppressed decay modes $\bar{B}_{s}^{0}\rightarrow\pi^{-}D^{+}$ and $\bar{B}_{s}^{0}\rightarrow\pi^{0}D^{0}$ are necessary, whose branching ratios we predict in Eq. (39).
[…]
@@
From our parameter extraction in Table 4 it is evident that many of the SU(3)${}_{F}$ breaking parameters are basically unconstrained. Future more precise branching ratio data would allow to test the pattern of the SU(3)${}_{F}$ anatomy with much more accuracy. Besides the unmeasured decay channels $\bar{B}_{s}^{0}\rightarrow\pi^{-}D^{+}$ and $\bar{B}_{s}^{0}\rightarrow\pi^{0}D^{0}$, there is a lot of room for improvement left in the channels $\bar{B}^{0}\rightarrow K^{-}D_{s}^{+}$, $\bar{B}_{s}^{0}\rightarrow K^{0}D^{0}$, $\bar{B}^{0}\rightarrow\bar{K}^{0}D^{0}$ and $\bar{B}_{s}^{0}\rightarrow K^{-}D_{s}^{+}$, all of which still have relative branching ratio uncertainties of $>10\%$. Being a general characteristic of symmetry-based methods, we need improvements in several decay channels in order to obtain a complete picture of the underlying dynamics of $B\rightarrow DP$ decays.
@@ Acknowledgements.
J.D. is supported by the European Research Council under the starting grant Beauty2Charm 852642. The work of A.S. is supported in part by the US DOE Contract No. DE-SC 0012704. S.S. is supported by a Stephen Hawking Fellowship from UKRI under reference EP/T01623X/1 and the STFC research grants ST/T001038/1 and ST/X00077X/1. For the purpose of open access, the authors have applied a Creative Commons Attribution (CC BY) licence to any Authors Accepted Manuscript version arising. This work uses existing data which is available at locations cited in the bibliography.

### v2

@@ IV Conclusions
Motivated by recently found discrepancies between experimental data and QCD factorization, we study the anatomy of a plain SU(3)F expansion in $B\rightarrow DP$ decays as extracted from the data, with no further theory assumptions. We find that, although the SU(3)F limit is excluded by $>5\sigma$, the data can be explained with amplitude-level SU(3)F breaking of $\sim 20\%$i.e., the SU(3)F expansion works to the expected level.
@@
As is well known, SU(3)F is an approximate symmetry of QCD as the light ($u$, $d$, $s$) quark masses are not equal to each other though they (especially $u$ and $d$) are small compared to $\Lambda_{\mathrm{QCD}}$. So when predictions made by assuming SU(3)F are not upheld experimentally, it often becomes important to estimate the size of SU(3)F breaking before invoking new physics.
@@
As one more piece in the puzzle, our results inform future theoretical and experimental work. By considering the complete system of decays related by the SU(3)F symmetry, we show unambiguously that the underlying issue is not connected to SU(3)F breaking, as suggested by the ratio in Eq. (1), which is in agreement with the data. In a scenario beyond the SM (BSM), any new physics model designed to explain the deviations has therefore to respect the SU(3)F structure of the SM operators, implying bounds on the hierarchy between BSM couplings to the first and second generations of quarks.
@@
It would be very interesting to complete the picture in the future by testing the isospin sum rule between $\mathcal{B}(\bar{B}_{s}^{0}\rightarrow\pi^{-}D^{+})$ and $\mathcal{B}(\bar{B}_{s}^{0}\rightarrow\pi^{0}D^{0})$, as well as the SU(3)F sum rule between $\mathcal{B}(\bar{B}^{0}\rightarrow K^{-}D_{s}^{+})$ and $\mathcal{B}(\bar{B}_{s}^{0}\rightarrow\pi^{-}D^{+})$, see Eqs. (23, 24, 29). To achieve this, measurements of the suppressed decay modes $\bar{B}_{s}^{0}\rightarrow\pi^{-}D^{+}$ and $\bar{B}_{s}^{0}\rightarrow\pi^{0}D^{0}$ are necessary, whose branching ratios we predict in Eq. (40).
The LHCb Upgrade II experiment has strong prospects for these modes. Run 5 of the LHC, scheduled for 2035, will see LHCb operating at instantaneous luminosities an order of magnitude greater than previously and accumulating a data sample corresponding to a minimum of 300 fb-1. This integrated luminosity, accompanied by improvements to the electromagnetic calorimeter granularity and energy resolution, will provide unprecedented sensitivity to modes with neutral particles in the final state Aaij et al. (2018).
@@
From our parameter extraction in Table 4 it is evident that many of the SU(3)F breaking parameters are basically unconstrained. Future more precise branching ratio data would allow one to test the pattern of the SU(3)F anatomy with much more accuracy. Besides the unmeasured decay channels $\bar{B}_{s}^{0}\rightarrow\pi^{-}D^{+}$ and $\bar{B}_{s}^{0}\rightarrow\pi^{0}D^{0}$, there is a lot of room for improvement left in the channels $\bar{B}^{0}\rightarrow K^{-}D_{s}^{+}$, $\bar{B}_{s}^{0}\rightarrow K^{0}D^{0}$, $\bar{B}^{0}\rightarrow\bar{K}^{0}D^{0}$ and $\bar{B}_{s}^{0}\rightarrow K^{-}D_{s}^{+}$, all of which still have relative branching ratio uncertainties of $>10\%$. Being a general characteristic of symmetry-based methods, we need improvements in several decay channels in order to obtain a complete picture of the underlying dynamics of $B\rightarrow DP$ decays.
@@ Acknowledgements.
J.D. is supported by the European Research Council under the starting grant Beauty2Charm 852642. The work of A.S. is supported in part by the US DOE Contract No. DE-SC 0012704. S.S. is supported by a Stephen Hawking Fellowship from UKRI under reference EP/T01623X/1 and the STFC research grants ST/T001038/1 and ST/X00077X/1.